# Topological characterization of phase transitions and critical edge states in one-dimensional non-Hermitian systems with sublattice symmetry

### v1

@@ III.3 $(\alpha,\alpha^{\prime})$ chain: topologically trivial/nontrivial critical points
The theoretical development in Sec. II allows us to identify $f(z)=J_{\alpha}^{{\rm R}}z^{-\alpha}+J_{\alpha^{\prime}}^{{\rm R}}z^{-\alpha^{\prime}}$ and $g(z)=J_{\alpha}^{{\rm L}}z^{\alpha}+J_{\alpha^{\prime}}^{{\rm L}}z^{\alpha^{\prime}}$ for the $(\alpha,\alpha^{\prime})$ chain. The characteristic functions $P(z)$ and $Q(z)$ are thus given by
$P(z)$ $=\frac{J_{\alpha}^{{\rm L}}z^{\alpha}+J_{\alpha^{\prime}}^{{\rm L}}z^{\alpha^{\prime}}}{J_{\alpha}^{{\rm R}}z^{-\alpha}+J_{\alpha^{\prime}}^{{\rm R}}z^{-\alpha^{\prime}}},$ (15)
@@
The GBZ of NHSSH $(\alpha,\alpha^{\prime})$ chain can be obtained from the two middle solutions of the equation $Q(z)=Q(ze^{i\theta})$ for $\theta\in[0,2\pi]$ [26]. Direct calculations yield the solution $z^{2(\alpha^{\prime}-\alpha)}=J_{\alpha}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}e^{-i(\alpha^{\prime}-\alpha)\theta}/(J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha}^{{\rm R}})$. Therefore, the GBZ of the $(\alpha,\alpha^{\prime})$ chain has the shape of a circle with the radius
$r_{0}=\left|\frac{J_{\alpha}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}}{J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha}^{{\rm R}}}\right|^{\frac{1}{2(\alpha^{\prime}-\alpha)}}.$ (17)
Under the non-Hermitian condition $J_{\alpha}^{{\rm L}}\neq(J_{\alpha}^{{\rm R}})^{*}$ and/or $J_{\alpha^{\prime}}^{{\rm L}}\neq(J_{\alpha^{\prime}}^{{\rm R}})^{*}$, we will generally have $r_{0}\neq 1$, making the GBZ different from the standard BZ of a Hermitian model that corresponds to a unit circle on the complex plane. The critical (gap-closing) points of the NHSSH $(\alpha,\alpha^{\prime})$ chain can be obtained by inserting the GBZ solution $z_{0}=r_{0}e^{i\phi}$ ($\phi\in[0,2\pi]$) into the equality $Q(z)=[E(z)]^{2}=0$, yielding the phase boundary equation in the hopping parameter space as
$|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|=|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|.$ (18)
The parameter regions with $|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|<|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|$ and $|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|>|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|$ are then expected to belong to distinct topological phases.
We next determine the locations of zeros/poles of $P(z)$ and $Q(z)$ relative to the GBZ. Using the relevant radius $r^{{\rm L}}=|J_{\alpha}^{{\rm L}}/J_{\alpha^{\prime}}^{{\rm L}}|^{1/(\alpha^{\prime}-\alpha)}$ and $r^{{\rm R}}=|J_{\alpha^{\prime}}^{{\rm R}}/J_{\alpha}^{{\rm R}}|^{1/(\alpha^{\prime}-\alpha)}$, we obtain
$\frac{r^{{\rm L}}}{r_{0}}=\left|\frac{J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}}{J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}}\right|^{\frac{1}{2(\alpha^{\prime}-\alpha)}},\quad\frac{r^{{\rm R}}}{r_{0}}=\left|\frac{J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}}{J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}}\right|^{\frac{1}{2(\alpha^{\prime}-\alpha)}}.$ (19)
Therefore, if $|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|<|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|$, the zeros/poles lying along the circles of radius $r^{{\rm L}}$ and $r^{{\rm R}}$ are outside and inside the GBZ, respectively. On the contrary, when $|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|>|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|$, the zeros/poles lying along the circles of radius $r^{{\rm L}}$ and $r^{{\rm R}}$ are inside and outside the GBZ, respectively. If $|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|=|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|$, we have $r^{{\rm L}}=r^{{\rm R}}=r_{0}$, which means that the zeros/poles along the circles of radius $r^{{\rm L}}$ and $r^{{\rm R}}$ are both on the GBZ when the phase boundary is reached. Finally, the zeros and poles at $z=0$ are all inside the GBZ so long as the GBZ radius $r_{0}>0$.
@@
$W_{1}=\begin{cases}2\alpha,&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|<|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\\ \alpha+\alpha^{\prime},&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|=|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\\ 2\alpha^{\prime},&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|>|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\end{cases},$ (20)
$W_{2}=\begin{cases}0,&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|<|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\\ \alpha-\alpha^{\prime},&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|=|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\\ 0,&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|>|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\end{cases}.$ (21)
Put together, we arrive at the following topological invariant and bulk-edge correspondence of the NHSSH $(\alpha,\alpha^{\prime})$ chain assuming $0\leq\alpha<\alpha^{\prime}$, i.e.,
$W=\frac{W_{1}+W_{2}}{2}=\begin{cases}\alpha,&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|\leq|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\\ \alpha^{\prime},&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|>|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\end{cases},$ (22)
$N_{0}=2|W|=\begin{cases}2\alpha,&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|\leq|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\\ 2\alpha^{\prime},&|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|>|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|\end{cases}.$ (23)
We conclude that the critical points along $|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|=|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|$ are topologically nontrivial whenever $\alpha>0$. Their nontrivial topologies are characterized by a quantized winding number $W=\alpha$ and a number $N_{0}=2\alpha$ of edge modes coexisting with a gapless bulk at zero energy. Instead, the critical points along $|J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}|=|J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}|$ are topologically trivial when $\alpha=0$, regardless of the value of $\alpha^{\prime}$. Away from the critical points, we find two distinct gapped phases with the winding numbers $\alpha,\alpha^{\prime}$ and the numbers of edge zero modes $2\alpha,2\alpha^{\prime}$. These gapped phases and gapless phase boundaries could thus be completely described within our theoretical framework.
@@
$r_{0}=\sqrt{|J-\gamma|/|J+\gamma|},\quad J^{2}=\gamma^{2}\pm 1.$ (24)
@@
These theoretical results can be verified by computing the zero energy solutions of the system explicitly. In the lattice representation, the Hamiltonian of NHSSH $(0,1)$ chain reads $\hat{H}_{01}=\hat{H}_{0}+\hat{H}_{1}$, where $\hat{H}_{0}=\sum_{n}\left[(J+\gamma)\hat{a}_{n}^{\dagger}\hat{b}_{n}+(J-\gamma)\hat{b}_{n}^{\dagger}\hat{a}_{n}\right]$ and $\hat{H}_{1}=\sum_{n}\left(\hat{b}_{n}^{\dagger}\hat{a}_{n+1}+\hat{a}_{n+1}^{\dagger}\hat{b}_{n}\right)$. For a chain with unit cell indices n=1,2,,Nn=1,2,...,N under the OBC, the (non-normalized) zero-energy eigenstates can be found in the limit $N\rightarrow\infty$ as (see Appendix A for details)
@@
where $|\varrho|=\sqrt{|J+\gamma|/|J-\gamma|}$. These states could describe a pair of edge zero modes if and only if $|J^{2}-\gamma^{2}|<1$ (see also Appendix A), which is exactly the condition for the system to be in its gapped topological phase with the winding number $W=1$ and the number of zero-energy edge modes $N_{0}=2$, as reported in Eqs. (25) and (26). Our theoretical predictions are thus confirmed. Notably, the edge zero modes $|\psi_{0}^{{\rm L}}\rangle$ and $|\psi_{0}^{{\rm R}}\rangle$ become delocalized at the critical points $J^{2}-\gamma^{2}=\pm 1$. The gapless phase boundary of the NHSSH $(0,1)$ chain is thus topologically trivial, which is characterized by the winding number $W=0$ and the number of edge zero modes $N_{0}=0$.
@@
In Fig. 5, we present the spectra and edge states of the NHSSH $(0,1)$ chain. In Fig. 5(a), the blue dots correspond to the spectrum obtained under the OBC, whereas the red line describes the spectrum gap $\min_{k\in[-\pi,\pi]}|E(k)|$ obtained under the PBC. It is clear that the PBC and OBC spectra have different gap closing points, which demonstrates the breakdown of Hermitian bulk-edge correspondence and the necessity of incorporating the non-Bloch band theory to characterize the NHSSH $(0,1)$ chain. In the parameter region $\gamma\in(1,\sqrt{3})$, we further observe eigenmodes pinned at zero energy, which represent topological edge modes. As illustrated by the OBC spectra in Figs. 5(b) and 5(c), these edge modes are absent at the critical points $\gamma=1,\sqrt{3}$, which confirms the topological triviality of gap closing (phase transition) points in the NHSSH $(0,1)$ chain. The OBC spectrum at $\gamma=1.3$ in Fig. 5(d) instead holds a pair of zero modes in the bulk gap, whose spatial profiles are shown in Fig. 5(e). These zero modes are indeed localized at the edges of the lattice, which verifies the prediction of winding number $W=1$ and number of edge modes $N_{0}=2$ at $(J,\gamma)=(\sqrt{2},1.3)$ in the phase diagram Fig. 4(a). Overall, the results presented in Fig. 5 are consistent with our theoretical descriptions, which further affirms the applicability of our approach to the topological characterization of both gapped and gapless phases in sublattice-symmetric non-Hermitian systems.
@@
The Eqs. (28) and (29) can also be verified by computing the zero-energy solutions of the system. In the lattice representation, the Hamiltonian of NHSSH $(1,2)$ chain takes the form $\hat{H}_{12}=\hat{H}_{1}+\hat{H}_{2}$, where $\hat{H}_{1}=\sum_{n}\left[(J-\gamma)\hat{b}_{n}^{\dagger}\hat{a}_{n+1}+(J+\gamma)\hat{a}_{n+1}^{\dagger}\hat{b}_{n}\right]$ and $\hat{H}_{2}=\sum_{n}\left(\hat{b}_{n}^{\dagger}\hat{a}_{n+2}+\hat{a}_{n+2}^{\dagger}\hat{b}_{n}\right)$. For a chain with unit cell indices n=1,2,,Nn=1,2,...,N under the OBC, the (non-normalized) zero-energy eigenstates are found in the limit $N\rightarrow\infty$ as (see Appendix A for details)
$|\varphi_{0}^{{\rm L,1}}\rangle=$ $\hat{a}_{1}^{\dagger}|\emptyset\rangle,$
@@
where $|\varrho|=\sqrt{|J+\gamma|/|J-\gamma|}$. The eigenmodes $|\varphi_{0}^{{\rm L,1}}\rangle$ and $|\varphi_{0}^{{\rm L,2}}\rangle$ persist at the two ends of the chain throughout the parameter space. They could thus survive at the critical points $J^{2}=\gamma^{2}\pm 1$, making the latter topologically nontrivial with $W=1$ and the number of edge zero modes $N_{0}=2$. Meanwhile, the states $|\varphi_{0}^{{\rm L,2}}\rangle$ and $|\varphi_{0}^{{\rm R,2}}\rangle$ form a pair of edge zero modes if and only if $|J^{2}-\gamma^{2}|<1$ (see Appendix A for details). Therefore, they could only survive in the gapped phase of the chain with $W=2$ and the number of edge zero modes $N_{0}=4$, as described by the Eqs. (28) and (29). In the region $|J^{2}-\gamma^{2}|>1$, we are left with the edge zero modes $|\varphi_{0}^{{\rm L,1}}\rangle$ and $|\varphi_{0}^{{\rm L,2}}\rangle$, which is coincident with the gapped topological phase of $W=1$ and $N_{0}=2$ according to the Eqs. (28) and (29). These observations confirm that our theory is valid in both the gapped topological phases and along the gapless phase boundaries of the NHSSH $(1,2)$ chain. Importantly, the nontrivial critical points of non-Bloch bands are correctly depicted within our topological characterization.
@@
In Fig. 7, we present the spectra and edge states of the NHSSH $(1,2)$ chain. We observe that under the OBC, zero energy eigenmodes persist throughout the spectrum (blue dots) in Fig. 7(a). Meanwhile, the gap closing points under the PBC and OBC are again different, making it necessary to employ the non-Bloch band theory for topological characterizations. Notably, we observe strongly localized zero modes in the spectra at the critical points $(J,\gamma)=(\sqrt{2},1)$ and $(\sqrt{2},\sqrt{3})$, as presented in Figs. 7(b) and 7(c). It implies that in stark contrast to the NHSSH $(0,1)$ chain, the band touching points of NHSSH $(1,2)$ chain are topologically nontrivial. Indeed, the zero modes in Figs. 7(b) and 7(c) are both localized at the two ends of the chain, forming a pair of degenerate edge modes at $E=0$ as shown in Fig. 7(d). On the other hand, our bulk theory yields the winding number $W=1$ at both critical points, leading to the consistent bulk-edge correspondence $N_{0}=2|W|$. Therefore, beyond previous approaches, our theory allows us to identify a unique class of topologically nontrivial gapless points between non-Bloch bands and characterize the associated bulk-edge correspondence in 1D non-Hermitian systems.
@@
For completeness, we present the spectra and edge states of the NHSSH $(1,2)$ chain for three gapped phases in Fig. 8, whose system parameters are given by $(J,\gamma)=(\sqrt{2},0.5)$, $(\sqrt{2},1.3)$ and $(\sqrt{2},2)$ in the phase diagram Fig. 6(a). According to the phase diagram, the topological invariants of these phases are $W=1,2,1$, while the numbers of edge zero modes in Figs. 8(d)–8(f) are $N_{0}=2,4,2$. The bulk-edge correspondence $N_{0}=2|W|$ is thus verified for these gapped phases. In conclusion, we have demonstrated the applicability of our theory to the topological characterization of NHSSH $(\alpha,\alpha^{\prime})$ chains for both gapped and gapless phases. In the next subsection, we study a more complicated class of NHSSH chain and reveal the presence of phase transitions along topological phase boundaries therein.
III.4 $(\alpha,\alpha^{\prime},\alpha^{\prime\prime})$ chain: critical topological phase transitions

### v2

@@ III.2 $(\alpha,\alpha^{\prime})$ chain: topologically trivial/nontrivial critical points
The theory outlined in Sec. II allows us to identify $f(z)=J_{\alpha}^{{\rm R}}z^{-\alpha}+J_{\alpha^{\prime}}^{{\rm R}}z^{-\alpha^{\prime}}$ and $g(z)=J_{\alpha}^{{\rm L}}z^{\alpha}+J_{\alpha^{\prime}}^{{\rm L}}z^{\alpha^{\prime}}$ for the $(\alpha,\alpha^{\prime})$ chain. The characteristic functions $P(z)$ and $Q(z)$ are thus given by
$P(z)$ $=({J_{\alpha}^{{\rm L}}z^{\alpha}+J_{\alpha^{\prime}}^{{\rm L}}z^{\alpha^{\prime}}})/({J_{\alpha}^{{\rm R}}z^{-\alpha}+J_{\alpha^{\prime}}^{{\rm R}}z^{-\alpha^{\prime}}}),$ (10)
@@
The GBZ of NHSSH $(\alpha,\alpha^{\prime})$ chain can be obtained from the two middle solutions of the equation $Q(z)=Q(ze^{i\theta})$ for $\theta\in[0,2\pi]$ [26]. Direct calculations yield the solution $z^{2(\alpha^{\prime}-\alpha)}=J_{\alpha}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}e^{-i(\alpha^{\prime}-\alpha)\theta}/(J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha}^{{\rm R}})$. Therefore, the GBZ of $(\alpha,\alpha^{\prime})$ chain has a circular shape with radius
$r_{0}=\left|{J_{\alpha}^{{\rm L}}J_{\alpha^{\prime}}^{{\rm R}}}/({J_{\alpha^{\prime}}^{{\rm L}}J_{\alpha}^{{\rm R}}})\right|^{\frac{1}{2(\alpha^{\prime}-\alpha)}}.$ (12)
Under the non-Hermitian condition $J_{\alpha}^{{\rm L}}\neq(J_{\alpha}^{{\rm R}})^{*}$ and/or $J_{\alpha^{\prime}}^{{\rm L}}\neq(J_{\alpha^{\prime}}^{{\rm R}})^{*}$, we generally have $r_{0}\neq 1$, making the GBZ different from the standard BZ of a Hermitian model that corresponds to a unit circle on the complex plane. The critical (gap-closing) points of the NHSSH $(\alpha,\alpha^{\prime})$ chain can be obtained by inserting the GBZ solution $z_{0}=r_{0}e^{i\phi}$ ($\phi\in[0,2\pi]$) into the equality $Q(z)=[E(z)]^{2}=0$, yielding the phase boundary in the hopping parameter space
$|{\cal J}_{\alpha^{\prime}}|=|{\cal J}_{\alpha}|,\qquad{\cal J}_{\alpha}\equiv{J_{\alpha}^{{\rm L}}J_{\alpha}^{{\rm R}}}\quad\forall\alpha.$ (13)
The parameter regions $|{\cal J}_{\alpha^{\prime}}|<|{\cal J}_{\alpha}|$ and $|{\cal J}_{\alpha^{\prime}}|>|{\cal J}_{\alpha}|$ are then expected to belong to distinct topological phases.
We next determine the locations of zeros/poles of $P(z)$ and $Q(z)$ relative to the GBZ. Using the relevant radii $r^{{\rm L}}$ and $r^{{\rm R}}$, we obtain
$\frac{r^{{\rm L}}}{r_{0}}=\left|{{\cal J}_{\alpha}}/{{\cal J}_{\alpha^{\prime}}}\right|^{\frac{1}{2(\alpha^{\prime}-\alpha)}},\quad\frac{r^{{\rm R}}}{r_{0}}=\left|{{\cal J}_{\alpha^{\prime}}}/{{\cal J}_{\alpha}}\right|^{\frac{1}{2(\alpha^{\prime}-\alpha)}}.$ (14)
Therefore, if $|{{\cal J}_{\alpha^{\prime}}}|<|{{\cal J}_{\alpha}}|$, the zeros/poles lying along the circles of radii $r^{{\rm L}}$ and $r^{{\rm R}}$ are outside and inside the GBZ, respectively. On the contrary, when $|{{\cal J}_{\alpha^{\prime}}}|>|{{\cal J}_{\alpha}}|$, the zeros/poles lying along the circles of radii $r^{{\rm L}}$ and $r^{{\rm R}}$ are inside and outside the GBZ, respectively. If $|{{\cal J}_{\alpha^{\prime}}}|=|{{\cal J}_{\alpha}}|$, we have $r^{{\rm L}}=r^{{\rm R}}=r_{0}$, which means that the zeros/poles along the circles of radii $r^{{\rm L}}$ and $r^{{\rm R}}$ are both on the GBZ when the phase boundary is reached. Finally, the zeros and poles at $z=0$ are all inside the GBZ so long as the GBZ radius $r_{0}>0$.
@@
$(W_{1},W_{2})=\begin{cases}(2\alpha,0),&|{{\cal J}_{\alpha^{\prime}}}|<|{{\cal J}_{\alpha}}|\\ (\alpha+\alpha^{\prime},\alpha-\alpha^{\prime}),&|{{\cal J}_{\alpha^{\prime}}}|=|{{\cal J}_{\alpha}}|\\ (2\alpha^{\prime},0),&|{{\cal J}_{\alpha^{\prime}}}|>|{{\cal J}_{\alpha}}|\end{cases},$ (15)
Put together, we arrive at the following topological invariant and bulk-edge correspondence of the NHSSH $(\alpha,\alpha^{\prime})$ chain assuming $0\leq\alpha<\alpha^{\prime}$ for any $\alpha,\alpha^{\prime}$, i.e.,
$W$ $=\frac{W_{1}+W_{2}}{2}=\begin{cases}\alpha,&|{\cal J}_{\alpha^{\prime}}|\leq|{\cal J}_{\alpha}|\\ \alpha^{\prime},&|{\cal J}_{\alpha^{\prime}}|>|{\cal J}_{\alpha}|\end{cases},$ (16)
$N_{0}$ $=2|W|=\begin{cases}2\alpha,&|{\cal J}_{\alpha^{\prime}}|\leq|{\cal J}_{\alpha}|\\ 2\alpha^{\prime},&|{\cal J}_{\alpha^{\prime}}|>|{\cal J}_{\alpha}|\end{cases}.$ (17)
We conclude that the critical points along $|{\cal J}_{\alpha^{\prime}}|=|{\cal J}_{\alpha}|$ are topologically nontrivial whenever $\alpha>0$. Their nontrivial topologies are characterized by a quantized winding number $W=\alpha$ and a number $N_{0}=2\alpha$ of edge modes coexisting with a gapless bulk at zero energy. Instead, the critical points along $|{\cal J}_{\alpha^{\prime}}|=|{\cal J}_{\alpha}|$ are topologically trivial when $\alpha=0$, regardless of the value of $\alpha^{\prime}$. Away from the critical points, we find two distinct gapped phases with the winding numbers $\alpha,\alpha^{\prime}$ and the numbers of edge zero modes $2\alpha,2\alpha^{\prime}$. These gapped phases and gapless phase boundaries could thus be completely described within our theoretical framework.
@@
$r_{0}=\sqrt{|J-\gamma|/|J+\gamma|},\qquad J^{2}=\gamma^{2}\pm 1.$ (18)
@@
The Eqs. (18)–(20) can be verified by computing the zero energy solutions of the system explicitly. In lattice representation, the Hamiltonian of NHSSH $(0,1)$ chain reads $\hat{H}_{01}=\hat{H}_{0}+\hat{H}_{1}$, where $\hat{H}_{0}=\sum_{n}[(J+\gamma)\hat{a}_{n}^{\dagger}\hat{b}_{n}+(J-\gamma)\hat{b}_{n}^{\dagger}\hat{a}_{n}]$ and $\hat{H}_{1}=\sum_{n}(\hat{b}_{n}^{\dagger}\hat{a}_{n+1}+\hat{a}_{n+1}^{\dagger}\hat{b}_{n})$. For a chain with unit cell indices n=1,2,,Nn=1,2,...,N under the open boundary condition (OBC), the (non-normalized) zero-energy eigenstates can be found in the limit $N\rightarrow\infty$ in symmetrized basis as (see Appendix C for details)
@@
where $|\varrho|=\sqrt{|J+\gamma|/|J-\gamma|}$. These states could describe a pair of edge zero modes if and only if $|J^{2}-\gamma^{2}|<1$ (see also Appendix C), which is exactly the condition for the system to be in its gapped topological phase with the winding number $W=1$ and the number of zero-energy edge modes $N_{0}=2$, as reported in Eqs. (19) and (20). Our theoretical predictions are thus confirmed. Notably, the edge zero modes $|\psi_{0}^{{\rm L}}\rangle$ and $|\psi_{0}^{{\rm R}}\rangle$ become delocalized at the critical points $J^{2}-\gamma^{2}=\pm 1$. The gapless phase boundary of the NHSSH $(0,1)$ chain is thus topologically trivial, which is characterized by the winding number $W=0$ and the number of edge zero modes $N_{0}=0$.
@@
In Fig. 3, we present the spectra and edge states of the NHSSH $(0,1)$ chain. In Fig. 3(a), the blue dots correspond to the spectrum obtained under the OBC, whereas the red line describes the spectrum gap $\min_{k\in[-\pi,\pi]}|E(k)|$ obtained under the PBC. It is clear that the PBC and OBC spectra have different gap closing points, which demonstrates the breakdown of Hermitian bulk-edge correspondence and the necessity of incorporating the non-Bloch band theory to characterize the NHSSH $(0,1)$ chain. In the parameter region $\gamma\in(1,\sqrt{3})$, we further observe eigenmodes pinned at zero energy, which represent topological edge modes. As illustrated by the OBC spectra in Figs. 3(b) and 3(c), these edge modes are absent at the critical points $\gamma=1,\sqrt{3}$, which confirms the topological triviality of gap closing (phase transition) points in the NHSSH $(0,1)$ chain. The OBC spectrum at $\gamma=1.3$ in Fig. 3(d) instead holds a pair of zero modes in the bulk gap, whose spatial profiles are shown in Fig. 3(e). These zero modes are indeed localized at the edges of the lattice, which verifies the prediction of winding number $W=1$ and number of edge modes $N_{0}=2$ at $(J,\gamma)=(\sqrt{2},1.3)$ in the phase diagram Fig. 2(a). Overall, the results presented in Fig. 3 are consistent with our theoretical descriptions, which further affirms the applicability of our approach to the topological characterization of both gapped and gapless phases in sublattice-symmetric non-Hermitian systems.
@@
The Eqs. (22) and (23) can also be verified by computing the zero-energy solutions of the system. In the lattice representation, the Hamiltonian of NHSSH $(1,2)$ chain takes the form $\hat{H}_{12}=\hat{H}_{1}+\hat{H}_{2}$, where $\hat{H}_{1}=\sum_{n}[(J-\gamma)\hat{b}_{n}^{\dagger}\hat{a}_{n+1}+(J+\gamma)\hat{a}_{n+1}^{\dagger}\hat{b}_{n}]$ and $\hat{H}_{2}=\sum_{n}(\hat{b}_{n}^{\dagger}\hat{a}_{n+2}+\hat{a}_{n+2}^{\dagger}\hat{b}_{n})$. For a chain with unit cell indices n=1,2,,Nn=1,2,...,N under the OBC, the (non-normalized) zero-energy eigenstates are found in the limit $N\rightarrow\infty$ and in symmetrized basis as (see Appendix C for details)
$|\varphi_{0}^{{\rm L,1}}\rangle=$ $\hat{a}_{1}^{\dagger}|\emptyset\rangle,\qquad|\varphi_{0}^{{\rm R,1}}\rangle=\hat{b}_{N}^{\dagger}|\emptyset\rangle,$
@@
where $|\varrho|=\sqrt{|J+\gamma|/|J-\gamma|}$. The eigenmodes $|\varphi_{0}^{{\rm L,1}}\rangle$ and $|\varphi_{0}^{{\rm R,1}}\rangle$ persist at the two ends of the chain throughout the parameter space. They could thus survive at the critical points $J^{2}=\gamma^{2}\pm 1$, making the latter topologically nontrivial with $W=1$ and the number of edge zero modes $N_{0}=2$. Meanwhile, the states $|\varphi_{0}^{{\rm L,2}}\rangle$ and $|\varphi_{0}^{{\rm R,2}}\rangle$ form a pair of edge zero modes if and only if $|J^{2}-\gamma^{2}|<1$ (see Appendix C for details). Therefore, they could only survive in the gapped phase of the chain with $W=2$ and the number of edge zero modes $N_{0}=4$, as described by the Eqs. (22) and (23). In the region $|J^{2}-\gamma^{2}|>1$, we are left with the edge zero modes $|\varphi_{0}^{{\rm L,1}}\rangle$ and $|\varphi_{0}^{{\rm R,1}}\rangle$, which is coincident with the gapped topological phase of $W=1$ and $N_{0}=2$ according to the Eqs. (22) and (23). These observations confirm that our theory is valid in both the gapped topological phases and along the gapless phase boundaries of the NHSSH $(1,2)$ chain. Importantly, the nontrivial critical points of non-Bloch bands are correctly depicted within our topological characterization.
@@
In Fig. 5, we present the spectra and edge states of the NHSSH $(1,2)$ chain. We observe that under the OBC, zero energy eigenmodes persist throughout the spectrum (blue dots) in Fig. 5(a). Meanwhile, the gap closing points under the PBC and OBC are again different, making it necessary to employ the non-Bloch band theory for topological characterizations. Notably, we observe strongly localized zero modes in the spectra at the critical points $(J,\gamma)=(\sqrt{2},1)$ and $(\sqrt{2},\sqrt{3})$, as presented in Figs. 5(b) and 5(c). It implies that in stark contrast to the NHSSH $(0,1)$ chain, the band touching points of NHSSH $(1,2)$ chain are topologically nontrivial. Indeed, the zero modes in Figs. 5(b) and 5(c) are both localized at the two ends of the chain, forming a pair of degenerate edge modes at $E=0$ as shown in Fig. 5(d). On the other hand, our bulk theory yields the winding number $W=1$ at both critical points, leading to the consistent bulk-edge correspondence $N_{0}=2|W|$. Therefore, beyond previous approaches, our theory allows us to identify a unique class of topologically nontrivial gapless points between non-Bloch bands and characterize the associated bulk-edge correspondence in 1D non-Hermitian systems.
@@
For completeness, we present the spectra and edge states of the NHSSH $(1,2)$ chain for three gapped phases in Fig. 6, whose system parameters are given by $(J,\gamma)=(\sqrt{2},0.5)$, $(\sqrt{2},1.3)$ and $(\sqrt{2},2)$ in the phase diagram Fig. 4(a). According to the phase diagram, the topological invariants of these phases are $W=1,2,1$, while the numbers of edge zero modes in Figs. 6(d)–6(f) are $N_{0}=2,4,2$. The bulk-edge correspondence $N_{0}=2|W|$ is thus verified for these gapped phases. In conclusion, we have demonstrated the applicability of our theory to the topological characterization of NHSSH $(\alpha,\alpha^{\prime})$ chains for both gapped and gapless phases. In the next subsection, we study a more complicated class of NHSSH chain and reveal the presence of phase transitions along topological phase boundaries.
III.3 $(\alpha,\alpha^{\prime},\alpha^{\prime\prime})$ chain: critical topological phase transitions